OpenStudy (anonymous):

Can you solve the following problems using a TI89"? Suppose you buy a piece of office equipmet for $18000. After 4 years you sell I for a scrap value of $4000. The equiptment is depreciated linearly over 4 years. Te rate of depreciation of the piece of equipment is

OpenStudy (anonymous):

if it is linear the answer is \[\frac{180000-40000}{4}\] no need for a TI 89

OpenStudy (anonymous):

thing don't really depreciate linearly, but that is what you need to compute for this question

OpenStudy (anonymous):

In my professors lecture he wrote out y= mx+b: v=c-rt does that have anything to do with my problem?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%2818000-4000%29%2F4 not really , it is the slope